A photographer at a baseball game is behind home p
Answer #1

For the cameras we know the expression:

1/s0 +1/si = 1/f.

Where s0 = object distance

si = image distance

f= focal length.

By doing some algebric

1/si = 1/f - 1/s0.

si = s0f/(s0-f)

This is the image distance. Hence from this equation we can easily concluded that for farther objects we need to move only lene farther away from the film or detector.